# Using Large Language Models to Tackle Fundamental Challenges in Graph Learning: A Comprehensive Survey

### v1

@@ 3.2.2 Structure-Imbalanced Graph Learning
(1) Traditional Methods The main techniques for addressing structural imbalance in graph data are structural enhancement or debiasing. Some methods are dedicated to dynamically adjusting the neighbor aggregation weights to weaken the over-dominance effect of high-centrality nodes. DegFairGNN [152] introduces a generalized degree fairness constraint that reallocates neighborhood attention weights between high- and low-degree nodes to correct structural biases in GNN message passing, alleviating the issue of central nodes overly dominating the process. In addition, KG-Mixup [153] analyzes degree bias in knowledge graph completion, proposing a degree-aware contrastive loss function to dynamically adjust geometric constraints in the entity embedding space for high- and low-degree entities. Another approach starts from the perspective of subgraph structure enhancement and improves the expressiveness of the tail structure by generating diversified substructures. SOLT-GNN [154] designs a size-aware hierarchical pooling strategy that balances the representation distribution of graph structures of different sizes through subgraph cropping and feature decoupling, addressing the long-tailed structural imbalance in graph classification. SAILOR [155] introduces a structural enhancement-based framework for tail node representation learning, enhancing the visibility of low-degree nodes in the global topology through k-hop subgraph expansion and adversarial edge generation strategies. There are also methods that alleviate the number and topological imbalance of graph structures from the perspective of global structure reconstruction and regularity constraints. HiRe [156] constructs a hierarchical relational meta-learning framework that employs a meta-path-guided negative sampling mechanism to balance the structural coverage density of head and tail relations in knowledge graphs. QTIAH-GNN [157] develops a heterogeneous graph neural network that is aware of both quantity and topological imbalance, addressing node number imbalance and topological connection pattern imbalance through meta-relation-specific neighbor sampling and topological entropy regularization.
@@
Although existing methods have alleviated the problem of graph structural imbalance to a certain extent, they still have limitations. They often rely on manually designed constraint rules, which leads to limited generalization ability and difficulty in adapting to complex and changing graph structural deviations. At the same time, these methods tend to ignore the semantic associations of the global topology, and usually adopt static adjustment strategies, which cannot dynamically adapt to the structural evolution patterns implicit in the graph. These limitations [16, 13] highlight the need for a new technique that can take into account both semantic understanding and dynamic adaptability to more comprehensively solve the problem of graph structural imbalance.
(2) Methods Based on LLMs Compared to traditional methods, LLM-based methods can achieve semantically driven structure completion and enhancement by generating context-aware virtual edges or subgraph structures to supplement the topological connections of tail nodes, as shown in Figure 8. Moreover, LLMs deeply integrate semantic understanding and structural analysis, identifying the underlying semantic causes of structural imbalance through causal reasoning (e.g., the logical absence of long-tail relationships in knowledge graphs) and generating logically constrained suggestions, rather than relying solely on statistical distribution adjustments.
@@
SATKGC [141] proposes a contrastive learning framework based on subgraph awareness, generating semantic embeddings aligned with local subgraph structures using LLMs. It also designs a structure-aware contrastive loss function to enhance the logical coherence of long-tail relationships in knowledge graphs. MPIKGC [142] integrates LLMs from three perspectives: semantic enhancement, structural correction, and logical reasoning. Through dynamic weight allocation and logical rule generation, it addresses the relationship distribution imbalance in knowledge graph completion tasks. GraphEdit [140] develops an LLM-based iterative graph optimization framework, utilizing a semantic generation-topology validation mechanism to dynamically correct edge connections (such as adding virtual edges for tail nodes or pruning noisy connections), achieving collaborative optimization of global structural balance and local semantic consistency. LLM4RGNN [143] distills the reasoning capabilities of GPT-4 into a local LLM to identify malicious edges and, in conjunction with a language-model-based edge predictor, recovers a robust graph structure.
LLMs can semantically enrich or complete graph structures, but they typically need the graph to be represented in a suitable textual form. The advantage is that they might introduce connections or structural patterns that are semantically plausible, not random, thereby improving minority structures in meaningful ways.
@@ 3.3 Cross-Domain Heterogeneity in Graphs
Table 3: Literature Summary of LLM-based Methods for Handling Cross-Domain Heterogeneity in Graphs.
Type Domain Reference (Method—Citation) Text-Attributed Graph Learning Textual Attribute Alignment TAPE[166], LLMRec[167], MINGLE[168], GHGRL[169] Graph Foundation Model OFA[162], UniGraph[170], BooG[171], Hyper-FM[172] Multimodal Attributed Graph Learning MLLM-based Multimodal Alignment LLMRec[167], MAGB[21] Graph-Enhanced Multimodal Alignment MMGL[173], GraphAdapter[174], TouchUp-G[175] UniGraph2[176] Structural Heterogeneous Graph Learning Topological Graph Textualization LLMtoGraph[177], NLGraph[178], Talk-like-a-Graph[179] GPT4Graph[180], GITA[181], LLM4Hypergraph[182] Attributed Graph Textualization GraphText[183], WalkLM[184], Path-LLM[185] InstructGLM[186], MuseGraph[187], Graph-LLM[188] Graph Token Learning GNP[189], GraphToken[190], GraphGPT[191] LLaGA[192], TEA-GLM[193], HiGPT[194]
@@ 3.3.1 Text-Attributed Graph Learning
(1) Traditional Methods for Textual Attribute Modeling To handle the heterogeneity in textual data, early data integration methods extract structured information from unstructured textual inputs, thereby reducing variability across sources and producing unified features for downstream tasks [198, 199, 200]. However, these approaches require manually design of data schemes and training extraction models tailored to specific applications. In contrast, statistical methods, such as Bag of Words and TF-IDF [201], have been introduced to automatically generate unified feature vectors from text without relying on domain-specific design. These feature vectors are often used as inputs for graph learning methods, such as GNNs [47, 48, 49], which further incorporate structural information from the graph to obtain more effective node representations for addressing downstream tasks. For instance, TADW [202] approximates DeepWalk [42] using matrix factorization, where the TF-IDF feature vectors of node textual attributes serve as the initial feature matrix. Paper2vec [203] utilizes learnable Word2vec [204] text embeddings as initial node features, which are then trained by predicting whether two nodes belong to the same neighborhood in the graph. While these early statistical methods or shallow models offer solutions for mapping heterogeneous textual attributes into a unified space, their limited expressiveness cannot effectively capture complex features in these textural attributes. Moreover, these methods adopt text embeddings as fixed initial node features rather than integrating them with structural learning, limiting the potential of leveraging graph topology to enhance the textual attributes representation learning [195].
Several advanced methods explore approaches for modeling textual attributes that can better integrate with graph structures, aiming to improve the representation learning of both textual attributes and graph structure data simultaneously. GraphFormers [195] proposes a GNN-nested architecture where the transformer-based textual embedding modules and the GNN modules are nested and trained together. Heterformer [205] introduces virtual neighbor tokens that capture information from both text-rich and textless neighbors, allowing the model to simultaneously consider textual semantic information and structural information to embed the textual node attributes. GIANT [206] leverages a self-supervised learning task called neighborhood prediction to generate graph-dependent text embeddings, which are then used in downstream node classification tasks. E2EG [207] extends GIANT [206] by introducing an end-to-end model that simultaneously learns node classification and neighborhood prediction, improving the node classification performance and computational efficiency. METAG [208] employs a single pretrained language model to learn multiplex text representations, utilizing relation prior tokens to capture relation-specific signals while maintaining parameter efficiency.
@@
A number of methods leverage the use of LLMs to generate unified textual descriptions, which can be fed into a trainable smaller language model (LM) to extract task-specific embeddings, as depicted in Figure 10. TAPE [166] leverages the powerful language comprehension capabilities of LLMs to predict the category of a node based on its textual attributes and generate explanations for the prediction, which can be regarded as enhanced and aligned textual outputs. These aligned textual outputs are fed into a smaller language model to generate node feature vectors. As shown in [165], the domain shift is reduced in the LLM-enhanced text compared to the original textual attributes. Similarly, LLMRec [167] leverages LLMs to summarize user profiles and item attributes within a recommendation network, reducing the heterogeneity of the original attributes and producing aligned textual representations. These representations are then reprocessed by LLMs to generate unified embedding vectors for both users and items. Instead of generating unified textual descriptions, MINGLE [168] utilizes LLMs to map clinical notes and medical codes into a unified embedding space and uses these unified embedding vectors as initial node features for training a hypergraph neural network. GHGRL [169] utilizes LLMs to unify textual attributes on heterogeneous graphs based on format types and content types. Specifically, it first summarizes types information of all nodes using LLMs, which are then used to generate attribute summaries for each node based on the predicted types. These generated attribute summaries are then fed into a language model to generate unified feature vectors for training a graph neural network.
@@
By transforming categorical and numerical features into textual attributes with language-based descriptions, LLMs can extend their ability to handle the heterogeneity in these structured attributes, which do not naturally belong to the textual modality. For example, in molecular graph modeling, node attributes such as atom types or properties are typically represented as categorical or numerical values. By converting these attributes into descriptive text and even enriching with domain-specific explanations, LLMs can effectively understand their semantic meaning and unify them within a common representation space. Both OFA [162] and UniGraph [170] address attribute heterogeneity by constructing Text-Attributed Graphs (TAGs). These models enhance textual attributes by incorporating additional semantically rich contextual descriptions. Furthermore, they leverage domain knowledge to enrich non-textual attributes with textual representations, enabling LLMs to process diverse attributes in a unified manner and generate consistent embedding vectors. BooG [171] follows a similar approach but further employs a contrastive learning-based pre-training objective, which enhances the ability to learn expressive representations and generalize across different domains and downstream tasks. Similarly, Hyper-FM [172] leverages a language model to extract semantic features from cross-domain textual attributes on hypergraphs and integrates structural information through hierarchical high-order neighbor prediction.

### v2

@@ 3.2.2 Structure-Imbalanced Graph Learning
(1) Traditional Methods The main techniques for addressing structural imbalance in graph data can be broadly divided into two lines: debiasing-based approaches and structural enhancement-based approaches. From the debiasing perspective, several methods focus on dynamically adjusting neighbor aggregation weights to reduce the over-dominance of high-centrality nodes. For example, DegFairGNN [203] introduces a generalized degree fairness constraint that reallocates neighborhood attention weights between high- and low-degree nodes, thereby mitigating bias in GNN message passing. Extending this idea to knowledge graphs, KG-Mixup [204] analyzes degree bias in entity embeddings and proposes a degree-aware contrastive loss to balance geometric constraints between high- and low-degree entities. Together, these approaches highlight the importance of reweighting meachanisms for correcting degree-induced biases. On the other hand, structural enhancement approaches aim to enrich the representation capacity of underrepresented nodes or substructures. SOLT-GNN [205] improves graph classification by using a size-aware hierarchical pooling strategy that balances representation distributions through subgraph cropping and feature decoupling. Building on the idea of subgraph augmentation, SAILOR [206] strengthens the visibility of low-degree nodes in the global topology via $k$-hop subgraph expansion and adversarial edge generation. Both methods demonstrate that explicitly enhancing substructures can effectively alleviate long-tailed imbalance in graphs. A complementary line of work tackles imbalance from the perspective of global structure reconstruction and regularization. HiRe [207] develops a hierarchical relational meta-learning framework with a meta-path-guided negative sampling mechanism, balancing the structural coverage density between head and tail relations in knowledge graphs. Similarly, QTIAH-GNN [208] introduces a heterogeneous GNN that jointly considers quantity imbalance and topological imbalance, using meta-relation-specific neighbor sampling together with topological entropy regularization. These methods illustrate how global structural modeling can address both numerical and topological imbalance.
@@
Although existing methods have alleviated the problem of graph structural imbalance to a certain extent, they still have limitations. They often rely on manually designed constraint rules, which leads to limited generalization ability and difficulty in adapting to complex and changing graph structural deviations. At the same time, these methods tend to ignore the semantic associations of the global topology, and usually adopt static adjustment strategies, which cannot dynamically adapt to the structural evolution patterns implicit in the graph [19, 16].
(2) Methods Based on LLMs LLM-based methods offer a new paradigm for addressing structural imbalance in graphs by generating semantically meaningful edits—such as context-aware virtual edges or subgraph structures—that enhance connectivity for long-tail nodes (Figure 8). Unlike traditional GNN-based approaches, which learn structural patterns primarily from in-graph statistical distributions, LLM-based frameworks integrate rich prior semantic knowledge obtained from large-scale pretraining with graph-specific structures.
@@
For example, SATKGC [188] constructs subgraph-aware embeddings and uses contrastive losses to reinforce coherence among tail relations. MPIKGC [190] integrates LLMs for semantic enhancement, structural correction, and logic-driven edge generation, effectively mitigating relational imbalance. LLM4RGNN [191] transfers GPT-4’s reasoning capacity to detect and recover from malicious structural edits in a lightweight graph refinement framework. Complementing these, GraphEdit [187] proposes an iterative generate-and-validate loop, where LLMs suggest candidate edges which are then filtered by topological validators to ensure semantic and structural consistency. These approaches demonstrate that, when graphs are represented in text-compatible forms, LLMs can inject new connections that are not arbitrary, but semantically aligned—offering principled structure augmentation for underrepresented classes.
3.2.3 Evaluation for Handling Imbalance in Graphs
We summarize the evaluation protocols for LLM-based methods addressing graph imbalance, covering key datasets, metrics, and downstream tasks (see Table 2). Commonly used datasets include Cora [81], Citeseer [85], PubMed [82], and ogbn-arxiv [83] for node classification under class imbalance, and FB15k-237, WN18RR [94, 124] for knowledge graph completion in low-resource settings. Domain-specific benchmarks such as threat-dataset [176] and Twitter-HetDrug [178] are also adopted for specialized applications like fraud detection and mental health analysis.
Standard evaluation metrics vary by task: Accuracy and Macro-F1 are widely used for classification, while MRR and Hits@N are reported for link prediction and knowledge graph completion. For recommendation and retrieval tasks, AUC, MRR, and nDCG are common. Downstream tasks include few-shot node classification, sentiment analysis, knowledge graph construction, and drug repositioning, demonstrating the adaptability of LLMs in balancing both label and structural distributions. These evaluations highlight the effectiveness of LLM-based methods in enhancing model performance on underrepresented classes and nodes.
@@ 3.3 Cross-Domain Heterogeneity in Graphs
Table 3: LLM-based methods for handling cross-domain heterogeneity in graphs, grouped by domains and tasks, with representative methods, datasets, metrics, and downstream tasks.
Domains Tasks Methods Typical Datasets Common Metrics Downstream Tasks Text- Attributed Graph Learning Textual Attribute Alignment TAPE [86] Cora [81], PubMed [82], Arxiv-2023 [86], ogbn-arxiv [83], ogbn-products [83] Accuracy Node Classification LLMRec [217] MovieLens [218], Netflix [219] Recall, NDCG, Precision Item Recommendation MINGLE [220] MIMIC-III [221], CRADLE [220] Accuracy, AUC, AUPR, F1 Node Classification GHGRL [222] IMDB [223], DBLP [223], ACM [223], Wiki-CS [90], IMDB-RIR [222], DBLP-RID [222] Macro-F1, Micro-F1 Node Classification Graph Foundation Model OFA [213] Cora [81], PubMed [82], ogbn-arxiv [83], Wiki-CS [90], MOLHIV [224], MOLPCBA [224], FB15K237 [94], WN18RR [124], ChEMBL [225] Accuracy, AUC, AUPR Node Classification, Link Prediction, Graph Classification UniGraph [226] Cora [81], PubMed [82], ogbn-arxiv [83], ogbn-products [83], Wiki-CS [90], FB15K237 [94], WN18RR [124], MOLHIV [224], MOLPCBA [224] AUC Node Classification, Link Prediction, Graph Classification BooG [227] Cora [81], PubMed [82], ogbn-arxiv [83], Wiki-CS [90], MOLHIV [224], MOLPCBA [224] AUC Node Classification, Graph Classification Hyper-FM [228] Cora-CA-Text [228], Cora-CC-Text [228], Pubmed-CA-Text [228], Pubmed-CC-Text [228], AminerText [228], Arxiv-Text [228], Movielens-Text [228], IMDB-Text [228], GoodBook-Text [228], PPI-Text [228] Accuracy Node Classification Multimodal Attributed Graph Learning MLLM-based Multimodal Alignment LLMRec [217] MovieLens [218], Netflix [219] Recall, NDCG, Precision Item Recommendation MAGB [24] Cora [81], Wiki-CS [90], Ele-Photo [168], Flickr [229], Movies [24], Toys [24], Grocery [24], Reddit-S [24], Reddit-M [24] Accuracy, F1 Node Classification Graph-Enhanced Multimodal Alignment MMGL [230] WikiWeb2M [231] BLEU-4, ROUGE-L, CIDEr Section Summarization GraphAdapter [232] ImageNet [233], StandfordCars [234], UCF101 [235], Caltech101 [236], Flowers102 [237], SUN397 [238], DTD [239], EuroSAT [240], FGVCAircraft [241], OxfordPets [242], Food101 [243] Accuracy Image Classification TouchUp-G [244] ogbn-arxiv [83], ogbn-products [83], Books [244], Amazon-CP [244] MRR, Hits@N, Accuracy link prediction, node classification UniGraph2 [245] Cora [81], PubMed [82], ogbn-arxiv [83], ogbn-papers100M [83], ogbn-products [83], Wiki-CS [90], FB15K237 [94], WN18RR [124], Amazon-Sports [246], Amazon-Cloth [246], Goodreads-LP [246], Goodreads-NC [246], Ele-Fashion [246], WikiWeb2M [231] Accuracy, BLEU-4, ROUGE-L, CIDEr Node Classification, Edge Classification, Section Summarization Structural Heterogeneous Graph Learning Topological Graph Textualization LLMtoGraph [42] synthetic graph data [42] Accuracy, Positive Response Ratio Node Classification, Path Finding, Pattern Matching NLGraph [92] NLGraph [92] Accuracy, Partial Credit Score, Relative Error Path Finding, Pattern Matching, Topological Sort Talk-like-a-Graph [247] GraphQA [247] Accuracy Link Prediction, Pattern Matching GPT4Graph [248] ogbn-arxiv [83], MOLHIV [224], MOLPCBA [224], MetaQA [249] Accuracy Node Classification, Graph Classification, Graph Query Language Generation GITA [250] GVLQA [250] Accuracy Link Prediction, Pattern Matching, Path Finding, Topological Sort LLM4-Hypergraph [251] LLM4Hypergraph [251] Accuracy Isomorphism Recognition, Structure Classification, Link Prediction, Path Finding
Table 3: LLM-based methods for handling cross-domain heterogeneity in graphs, grouped by domains and tasks, with representative methods, datasets, metrics, and downstream tasks. (continued from previous page)
Domains Tasks Methods Typical Datasets Common Metrics Downstream Tasks Structural Heterogeneous Graph Learning Attributed Graph Textualization GraphText [252] Cora [81], Citeseer [85], Texas [253], Wisconsin [253], Cornell [253] Accuracy Node Classification WalkLM [254] PubMed [82], MIMIC-III [221] Macro-F1, Micro-F1, AUC, MRR Node Classification, Link Prediction Path-LLM [255] Cora [81], Citeseer [85], PubMed [82], ogbn-arxiv [83] Macro-F1, Micro-F1, AUC, Accuracy Node Classification, Link Prediction InstructGLM [256] Cora [81], PubMed [82], ogbn-arxiv [83] Accuracy Node Classification, Link Prediction MuseGraph [257] Cora [81], ogbn-arxiv [83], MIMIC-III [221], AGENDA [258], WebNLG [259] Macro-F1, Micro-F1, Weighted-F1, BLEU-4, METEOR, ROUGE-L, CHRF++ Node Classification, Graph-to-Text Generation Graph-LLM [260] Cora [81], Citeseer [85], PubMed [82], ogbn-arxiv [83], ogbn-products [83] Accuracy Node Classification Graph Token Learning GNP [261] OBQA [262], ARC [263], PIQA [264], Riddle [265], PQA [266], BioASQ [267] Accuracy Question Answering GraphToken [268] GraphQA [247] Accuracy Link Prediction, Pattern Matching GraphGPT [269] Cora [81], PubMed [82], ogbn-arxiv [83] Accuracy, Macro-F1, AUC Node Classification, Link Prediction LLaGA [270] Cora [81], PubMed [82], ogbn-arxiv [83], ogbn-products [83] Accuracy Node Classification, Link Prediction TEA-GLM [271] Cora [81], PubMed [82], ogbn-arxiv [83], TAG benchmark [168] Accuracy, AUC Node Classification, Link Prediction HiGPT [272] IMDB [223], DBLP [223], ACM [223] Macro-F1, Micro-F1, AUC Node Classification
@@ 3.3.1 Text-Attributed Graph Learning
(1) Traditional Methods for Textual Attribute Modeling To handle the heterogeneity in textual data, early data integration methods extract structured information from unstructured textual inputs, thereby reducing variability across sources and producing unified features for downstream tasks [276, 277, 278]. However, these approaches require manual design of data schemes and training extraction models tailored to specific applications. In contrast, statistical methods, such as Bag of Words and TF-IDF [279], have been introduced to automatically generate unified feature vectors from text without relying on domain-specific design. These feature vectors are often used as inputs for graph learning methods, such as GNNs [55, 56, 57], which further incorporate structural information from the graph to obtain more effective node representations for addressing downstream tasks. For instance, TADW [280] approximates DeepWalk [47] using matrix factorization, where the TF-IDF feature vectors of node textual attributes serve as the initial feature matrix. Paper2vec [281] utilizes learnable Word2vec [282] text embeddings as initial node features, which are then trained by predicting whether two nodes belong to the same neighborhood in the graph. While these early statistical methods or shallow models offer solutions for mapping heterogeneous textual attributes into a unified space, their limited expressiveness cannot effectively capture complex features in these textual attributes. Moreover, these methods adopt text embeddings as fixed initial node features rather than integrating them with structural learning, limiting the potential of leveraging graph topology to enhance the textual attributes representation learning [273].
Several advanced methods explore approaches for modeling textual attributes that can better integrate with graph structures, aiming to improve the representation learning of both textual attributes and graph structure data simultaneously. Some approaches, like GIANT [283] and its extension E2EG [284], leverage self-supervised learning to force text embeddings to encode graph-dependent information, which is trained by predicting a node’s neighbors from its text. Other methods focus on creating joint architectures. For example, GraphFormers [273] proposes a GNN-nested architecture where the transformer-based textual embedding modules and the GNN modules are nested and trained together. Heterformer [285] introduces virtual neighbor tokens that capture information from both text-rich and textless neighbors, allowing the model to simultaneously consider textual semantic information and structural information to embed the textual node attributes. To capture relation-specific signals in heterogeneous graphs, METAG [286] employs a single pretrained language model to learn multiplex text representations, which can effectively handle the diverse semantic relations while maintaining high parameter efficiency.
@@
A number of methods leverage the use of LLMs to generate unified textual descriptions, which can be fed into a trainable smaller language model to extract task-specific embeddings, as depicted in Figure 10. TAPE [86] leverages the powerful language comprehension capabilities of LLMs to predict the category of a node based on its textual attributes and generate explanations for the prediction, which can be regarded as enhanced and aligned textual outputs. These aligned textual outputs are fed into a smaller language model to generate node feature vectors. As shown in [216], the domain shift is reduced in the LLM-enhanced text compared to the original textual attributes. Similarly, LLMRec [217] leverages LLMs to summarize user profiles and item attributes within a recommendation network, reducing the heterogeneity of the original attributes and producing aligned textual representations. These representations are then reprocessed by LLMs to generate unified embedding vectors for both users and items. Instead of generating unified textual descriptions, MINGLE [220] utilizes LLMs to map clinical notes and medical codes into a unified embedding space and uses these unified embedding vectors as initial node features for training a hypergraph neural network. GHGRL [222] utilizes LLMs to unify textual attributes on heterogeneous graphs based on format types and content types. Specifically, it first summarizes types information of all nodes using LLMs, which are then used to generate attribute summaries for each node based on the predicted types. These generated attribute summaries are then fed into a language model to generate unified feature vectors for training a graph neural network.
@@
By transforming categorical and numerical features into textual attributes with language-based descriptions, LLMs can extend their ability to handle the heterogeneity in these structured attributes, which do not naturally belong to the textual modality. For example, in molecular graph modeling, node attributes such as atom types or properties are typically represented as categorical or numerical values. By converting these attributes into descriptive text and enriching with domain-specific explanations, LLMs can effectively understand their semantic meaning and unify them within a common representation space. Both OFA [213] and UniGraph [226] address attribute heterogeneity by constructing text-attributed graphs. These models enhance textual attributes by incorporating additional semantically rich contextual descriptions. Furthermore, they leverage domain knowledge to enrich non-textual attributes with textual representations, enabling LLMs to process diverse attributes in a unified manner and generate consistent embedding vectors. BooG [227] follows a similar approach but further employs a contrastive learning-based pretraining objective, which enhances the ability to learn expressive representations and generalize across different domains and downstream tasks. Similarly, Hyper-FM [228] leverages a language model to extract semantic features from cross-domain textual attributes on hypergraphs and integrates structural information through hierarchical high-order neighbor prediction.